# Drift estimation for a partially observed mixed fractional Ornstein–Uhlenbeck process

### v1

@@ 1 Introduction
$\mathcal{I}(\vartheta)=\frac{1}{2\vartheta}-\frac{2\vartheta}{\alpha(\alpha+\vartheta)}+\frac{\vartheta^{2}}{2\alpha^{3}},\,\alpha=\sqrt{\mu^{2}+\vartheta^{2}}.$
The rest of the paper is organised as follows. In Section 2 we recall the canonical innovation representation of the mixed fractional Brownian motion and construct the state–observation model corresponding to (1.1). Section 3 contains the Kalman–Bucy filtering equations and the Riccati equation for the error covariance. In Section 4 we introduce the Laplace transform of the squared distance between filters and derive its representation via a matrix Riccati equation. In Section 5 we analyse the long-time behaviour of the error covariance and prove the mixed Laplace condition $(L_{\mathrm{mix}})$; for clarity the argument is presented in six steps (Step 1Step 6). Section 6 identifies the Fisher information in closed form and shows that it coincides with the classical Kalman–Bucy expression.
@@ 4.1 Likelihood ratio and Laplace transform
where $1/8$ will be the same as in [1] with $\lambda=1/2$. From the likelihood function we can construct almost the same likelihood ratio and verify the same condition of (A.1)–(A.3) in [1]. also that is to say we will verify the same condition (L) for the Laplace transform of the quadratic form of the difference of the differnce
@@ 5.1 Rescaled covariance and its limit
The argument is based on rewriting the Riccati equation (3.3) in physical time, rescaling the covariance with $\Delta(t)$ and applying perturbation theory to the limiting autonomous Riccati equation.
Step 1. Riccati equation in physical time
From (3.3) and $\mathrm{d}\langle N\rangle_{t}/\mathrm{d}t=\psi(t)^{-1}$ we obtain for $\Gamma_{t}:=\gamma_{\zeta,\zeta}^{\vartheta}(t)$
@@
where
$F(t,X):=-\frac{\vartheta}{2}\bigl(A(t)X+XA(t)^{\top}\bigr)+b(t)b(t)^{\top}-\frac{\mu^{2}}{4}\,X\,\ell(t)\ell(t)^{\top}X.$ (5.3)
Here $A(t),b(t),\ell(t)$ are given by (2.7).
Step 2. Equation for the rescaled covariance
Set
$\widetilde{\Gamma}_{t}:=\widetilde{\gamma}^{\vartheta}(t)=\Delta(t)\Gamma_{t}\Delta(t).$
@@
$\dot{\widetilde{\Gamma}}_{t}=R(t)\widetilde{\Gamma}_{t}+\widetilde{\Gamma}_{t}R(t)^{\top}+\Delta(t)\dot{\Gamma}_{t}\Delta(t),$
where
$R(t):=\dot{\Delta}(t)\Delta(t)^{-1}.$
@@
Step 3. Bounds on $R(t)$
From (5.1) we have
@@
By (2.3), the scalar function $\dot{\psi}(t)/\psi(t)$ is square integrable on $t_{0},\infty)$ and tends to zero as $t\to\infty$. In particular,
$\lim_{t\to\infty}\|R(t)\|=0,\qquad\int_{t_{0}}^{\infty}\|R(t)\|\,\mathrm{d}t<\infty.$
Step 4. Exact autonomous form of the rescaled Riccati map
With the scaling $\Delta(t)$ in (5.1) and the definition $\widetilde{\Gamma}_{t}=\Delta(t)\Gamma_{t}\Delta(t)$, the nonlinearity in (5.4) admits the exact, time-independent form
@@
Hence the rescaled equation can be written as
$\dot{\widetilde{\Gamma}}_{t}=R(t)\widetilde{\Gamma}_{t}+\widetilde{\Gamma}_{t}R(t)^{\top}+\widetilde{F}_{\infty}(\widetilde{\Gamma}_{t}),$
where $R(t)$ is as in Step 3 and satisfies $\|R(t)\|\to 0$ and $\int_{t_{0}}^{\infty}\!\|R(t)\|\,dt<\infty$.
Step 5. Boundedness of $\widetilde{\Gamma}_{t}$
Using Step 4, the rescaled covariance solves
$\dot{\widetilde{\Gamma}}_{t}=R(t)\widetilde{\Gamma}_{t}+\widetilde{\Gamma}_{t}R(t)^{\top}+\widetilde{F}_{\infty}(\widetilde{\Gamma}_{t}),$
with $J=\bigl(\begin{matrix}1&1\\[-1.99997pt] 1&1\end{matrix}\bigr)$ and $\widetilde{F}_{\infty}(X)=-\tfrac{\vartheta}{2}(JX+XJ)+J-\tfrac{\mu^{2}}{4}XJX$. Let $u=\tfrac{1}{\sqrt{2}}(1,1)^{\top}$ and $v=\tfrac{1}{\sqrt{2}}(1,-1)^{\top}$ so that $J=2uu^{\top}$. In the $\{u,v\}$ basis write $x_{uu}=u^{\top}\widetilde{\Gamma}_{t}u$, $x_{uv}=u^{\top}\widetilde{\Gamma}_{t}v$, $x_{vv}=v^{\top}\widetilde{\Gamma}_{t}v$. A direct computation using $J=2uu^{\top}$ yields the autonomous system (the terms with $R(t)$ are collected in remainders $\mathcal{R}_{\bullet}$):
@@
Define the Lyapunov function $V(\widetilde{\Gamma}_{t}):=x_{uu}+\lambda x_{uv}^{2}+\beta x_{vv}$ with fixed $\lambda,\beta>0$. From the above display we obtain
$\dot{V}\leq C_{0}-\kappa_{0}V+c\,\|R(t)\|\,(1+V),$
for suitable constants $C_{0},\kappa_{0}>0$ depending only on $(\vartheta,\mu,\lambda,\beta)$. Since $\int_{t_{0}}^{\infty}\|R(t)\|\,dt<\infty$, a Gronwall–Bellman argument with integrable perturbation yields $\sup_{t\geq t_{0}}V(\widetilde{\Gamma}_{t})<\infty$. Hence $\{\widetilde{\Gamma}_{t}:t\geq t_{0}\}$ is bounded in $\mathbb{S}_{+}^{2}$, the cone of $2\times 2$ symmetric non-negative definite matrices.
Step 6. Identification of the limit
Consider the autonomous Riccati ODE $\dot{X}=\widetilde{F}_{\infty}(X)$ on $\mathbb{S}_{+}^{2}$. Solving $\widetilde{F}_{\infty}(X_{\infty})=0$ with the ansatz $X_{\infty}=g(\vartheta)J$ gives the scalar quadratic $\mu^{2}g^{2}+2\vartheta g-1=0$ with the unique nonnegative root $g(\vartheta)=(\sqrt{\vartheta^{2}+\mu^{2}}-\vartheta)/\mu^{2}$. Together with the boundedness in Step 5 and the integrable perturbation in Step 3, standard asymptotically autonomous ODE arguments imply $\widetilde{\Gamma}_{t}\to X_{\infty}(\vartheta)$ as $t\to\infty$. We denote this limit by $\widetilde{\Gamma}_{\infty}(\vartheta)$.
@@ 6.1 Asymptotical formula of Laplace transform
$\lim_{T\rightarrow\infty}L_{T}^{\mathrm{mix}}(a,\vartheta_{1},\vartheta_{2})=\exp\left\{-\frac{1}{2}\int_{0}^{T}\mathrm{tr}\,\mathcal{A}_{\infty}^{\mathrm{mix}}(t)\,\mathrm{d}\langle N\rangle_{t}\right\}\bigl(\det\Xi_{1,\infty}(T)\bigr)^{-1/2},$ (6.1)
where
@@
$\frac{\mathrm{d}}{\mathrm{d}\langle N\rangle_{t}}\Xi_{2,\infty}(t)$ $=\Xi_{1,\infty}(t)B(t)B(t)^{\top}+\Xi_{2,\infty}(t)\mathcal{A}_{\infty}^{\mathrm{mix}}(t)^{\top},\qquad\Xi_{2}(0)=0.$ (6.3)
Kronecker/Hamiltonian embedding. Recall from (2.7) that
@@ 6.2 Perturbation analysis of eigenvalues
1. Expansion for $x_{3}(h)$ via Schur Complement[9]. Since $x_{3}$ is a perturbation of $\alpha_{2}$, and $\alpha_{2}$ is separated from $\vartheta_{1}$, we can isolate the $2\times 2$ block corresponding to indices $\{2,4\}$ (the observation subsystem). Using the Schur complement with respect to this block, the characteristic equation near $x\approx\alpha_{2}$ reduces to:
@@
2. Expansion for $x_{1}(h)$ via Feedback Cycle Analysis. The eigenvalue $x_{1}$ corresponds to the drift parameter $\vartheta_{1}$. Let $x=\vartheta_{1}-\varepsilon$, where $\varepsilon$ is a small correction. Instead of a full block decomposition, we identify the dominant terms in the Leibniz expansion of the determinant.

### v2

@@ 1 Introduction
$\mathcal{I}(\vartheta)=\frac{1}{2\vartheta}-\frac{2\vartheta}{\alpha(\alpha+\vartheta)}+\frac{\vartheta^{2}}{2\alpha^{3}},\,\alpha=\sqrt{\mu^{2}+\vartheta^{2}}.$ (1.3)
The rest of the paper is organised as follows. In Section 2 we recall the canonical innovation representation of the mixed fractional Brownian motion and construct the state–observation model corresponding to (1.1). Section 3 contains the Kalman–Bucy filtering equations and the Riccati equation for the error covariance. In Section 4 we introduce the Laplace transform of the squared distance between filters and derive its representation via a matrix Riccati equation. In Section 5 we analyse the long-time behaviour of the error covariance and prove the mixed Laplace condition $(L_{\mathrm{mix}})$. Section 6 identifies the Fisher information in closed form and shows that it coincides with the classical Kalman–Bucy expression.
@@ 4.1 Likelihood ratio and Laplace transform
where $1/8$ will be the same as in [1] with $\lambda=1/2$. From the likelihood function we can construct almost the same likelihood ratio and verify the same condition of (A.1)–(A.3) in [1]. also that is to say we will verify the same condition (L) of [1] for the Laplace transform of the quadratic form of the difference of the differnce
@@
and in the following part we will try to achieve this limit and check that the fisher information $\mathcal{I}(\vartheta)$ satisfy the formula of (1.3).
@@ 5.1 Rescaled covariance and its limit
Proof.
The proof is based on rewriting the Riccati equation (3.3) in physical time, rescaling the covariance by $\Delta(t)$ from (5.1), and comparing the resulting equation with a limiting autonomous Riccati ODE.
First, we rewrite the Riccati equation in physical time. From (3.3) and $\mathrm{d}\langle N\rangle_{t}/\mathrm{d}t=\psi(t)^{-1}$ we obtain for $\Gamma_{t}:=\gamma_{\zeta,\zeta}^{\vartheta}(t)$,
@@
where $F$ is defined by
$F(t,X):=-\frac{\vartheta}{2}\bigl(A(t)X+XA(t)^{\top}\bigr)+b(t)b(t)^{\top}-\frac{\mu^{2}}{4}\,X\,\ell(t)\ell(t)^{\top}X,$
with $A(t),b(t),\ell(t)$ are given by (2.7).
Next, define the rescaled covariance
$\widetilde{\Gamma}_{t}:=\Delta(t)\Gamma_{t}\Delta(t).$
@@
$\dot{\widetilde{\Gamma}}_{t}=R(t)\widetilde{\Gamma}_{t}+\widetilde{\Gamma}_{t}R(t)^{\top}+\Delta(t)\dot{\Gamma}_{t}\Delta(t),\qquad R(t):=\dot{\Delta}(t)\Delta(t)^{-1}.$
@@
We now control the perturbation matrix $R(t)$. From (5.1),
@@
By (2.3), the scalar function $\dot{\psi}(t)/\psi(t)$ is square integrable on $t_{0},\infty)$ and tends to 0 as $t\to\infty$. In particular,
$\lim_{t\to\infty}\|R(t)\|=0,\qquad\int_{t_{0}}^{\infty}\|R(t)\|^{2}\,\mathrm{d}t<\infty.$
With the scaling (5.1), the nonlinearity in (5.3) admits the exact, time-independent form
@@
Hence (5.3) can be rewritten as
$\dot{\widetilde{\Gamma}}_{t}=R(t)\widetilde{\Gamma}_{t}+\widetilde{\Gamma}_{t}R(t)^{\top}+\widetilde{F}_{\infty}\bigl(\widetilde{\Gamma}_{t}\bigr).$ (5.4)
We next show that $\{\widetilde{\Gamma}_{t}:t\geq t_{0}\}$ is bounded in $\mathbb{S}_{+}^{2}$ without assuming $\int_{t_{0}}^{\infty}\|R(t)\|\,\mathrm{d}t<\infty$. Let $u=\tfrac{1}{\sqrt{2}}(1,1)^{\top}$ and $v=\tfrac{1}{\sqrt{2}}(1,-1)^{\top}$ so that $J=2uu^{\top}$. In the $\{u,v\}$ basis write
$x_{uu}=u^{\top}\widetilde{\Gamma}_{t}u,\qquad x_{uv}=u^{\top}\widetilde{\Gamma}_{t}v,\qquad x_{vv}=v^{\top}\widetilde{\Gamma}_{t}v.$
A direct computation (the terms with $R(t)$ are collected in remainders $\mathcal{R}_{\bullet}$) yields
@@
Define the Lyapunov function $V(\widetilde{\Gamma}_{t}):=x_{uu}+\lambda x_{uv}^{2}+\beta x_{vv}$ with fixed $\lambda,\beta>0$. Then there exist constants $C_{0},\kappa_{0}>0$ such that
$\dot{V}(t)\leq C_{0}-\kappa_{0}V(t)+c\,\|R(t)\|\,(1+V(t)),\qquad t\geq t_{0}.$
Since $\|R(t)\|\to 0$ as $t\to\infty$, there exists $T_{1}\geq t_{0}$ such that $c\|R(t)\|\leq\kappa_{0}/2$ for all $t\geq T_{1}$. Hence, for $t\geq T_{1}$,
$\dot{V}(t)\leq C_{0}+c\|R(t)\|-\frac{\kappa_{0}}{2}V(t)\leq C_{1}-\frac{\kappa_{0}}{2}V(t),$
where $C_{1}:=C_{0}+\sup_{t\geq T_{1}}c\|R(t)\|<\infty$. Solving this linear differential inequality yields
$V(t)\leq V(T_{1})e^{-\frac{\kappa_{0}}{2}(t-T_{1})}+\frac{2C_{1}}{\kappa_{0}},\qquad t\geq T_{1}.$
Therefore $\sup_{t\geq t_{0}}V(\widetilde{\Gamma}_{t})<\infty$, and consequently the trajectory $\{\widetilde{\Gamma}_{t}:t\geq t_{0}\}$ is bounded in $\mathbb{S}_{+}^{2}$.
Finally, we identify the limit by a fully expanded asymptotically autonomous argument. Consider the limiting autonomous Riccati ODE on $\mathbb{S}_{+}^{2}$,
$\dot{X}=\widetilde{F}_{\infty}(X).$
Solving $\widetilde{F}_{\infty}(X_{\infty})=0$ with the ansatz $X_{\infty}=g(\vartheta)J$ gives $\mu^{2}g^{2}+2\vartheta g-1=0$ and the unique nonnegative root
$g(\vartheta)=\frac{\sqrt{\vartheta^{2}+\mu^{2}}-\vartheta}{\mu^{2}},\qquad X_{\infty}(\vartheta)=g(\vartheta)J.$
[…]
We pass to the limit in the integral form of (5.4). For $s\in[-S,S]$,
$X_{n}(s)-X_{n}(0)=\int_{0}^{s}\Bigl(R(t_{n}+r)X_{n}(r)+X_{n}(r)R(t_{n}+r)^{\top}+\widetilde{F}_{\infty}(X_{n}(r))\Bigr)\,\mathrm{d}r.$
Because $\|R(t)\|\to 0$ as $t\to\infty$, we have $\sup_{|r|\leq S}\|R(t_{n}+r)\|\to 0$. Using uniform boundedness of $X_{n}$, uniform convergence $X_{n}\to X_{\ast}$, and continuity of $\widetilde{F}_{\infty}$, we can let $n\to\infty$ to obtain
$X_{\ast}(s)-X_{\ast}(0)=\int_{0}^{s}\widetilde{F}_{\infty}\bigl(X_{\ast}(r)\bigr)\,\mathrm{d}r,\qquad s\in[-S,S].$
[…]
Since the trajectory $\widetilde{\Gamma}_{t}$ is bounded and its $\omega$–limit set consists of equilibria, the limit $\lim_{t\to\infty}\widetilde{\Gamma}_{t}$ exists and equals an equilibrium. Denoting this limit by $\widetilde{\Gamma}_{\infty}(\vartheta)$, we conclude that $\widetilde{\Gamma}_{\infty}(\vartheta)=X_{\infty}(\vartheta)=g(\vartheta)J$, and hence (5.2) holds. ∎
@@ 6.1 Asymptotical formula of Laplace transform
$L_{T}^{\mathrm{mix}}(a,\vartheta_{1},\vartheta_{2})=\exp\left\{-\frac{1}{2}\int_{0}^{T}\mathrm{tr}\,\mathcal{A}_{\infty}^{\mathrm{mix}}(t)\,\mathrm{d}\langle N\rangle_{t}\right\}\bigl(\det\Xi_{1,\infty}(T)\bigr)^{-1/2}\,(1+o(1)),\qquad T\to\infty.$ (6.1)
where we replace the exact coefficients by their limits, resulting in a relative error $1+o(1)$ and here
@@
$\frac{\mathrm{d}}{\mathrm{d}\langle N\rangle_{t}}\Xi_{2,\infty}(t)$ $=\Xi_{1,\infty}(t)B(t)B(t)^{\top}+\Xi_{2,\infty}(t)\mathcal{A}_{\infty}^{\mathrm{mix}}(t)^{\top},\qquad\Xi_{2,\infty}(0)=0.$ (6.3)
Now, we are prepared to deal with the Kronecker or Hamiltonian embedding. Recall from (2.7) that
@@ 6.2 Perturbation analysis of eigenvalues
We first get the expansion for $x_{3}(h)$ via Schur Complement[9]. Since $x_{3}$ is a perturbation of $\alpha_{2}$, and $\alpha_{2}$ is separated from $\vartheta_{1}$, we can isolate the $2\times 2$ block corresponding to indices $\{2,4\}$ (the observation subsystem). Using the Schur complement with respect to this block, the characteristic equation near $x\approx\alpha_{2}$ reduces to:
@@
Then we analysis the expansion for $x_{1}(h)$ via Feedback Cycle Analysis. The eigenvalue $x_{1}$ corresponds to the drift parameter $\vartheta_{1}$. Let $x=\vartheta_{1}-\varepsilon$, where $\varepsilon$ is a small correction. Instead of a full block decomposition, we identify the dominant terms in the Leibniz expansion of the determinant.